How many ways are there to choose a valid password of the computer system
1
Expert's answer
2020-04-15T12:06:51-0400

A standard Qwerty keyboard can (by default) produce 96 unique characters.

Suppose we intend to create a password of size 'n'.

These each of the n places of this password can be filled using any of the above mentioned 96 characters.

Thus, total possible combinations would be "=96*96*96*....*96=96^n"

Thus, there are "96^n" ways to choose a valid password (of size 'n') for the computer system.